OpenStudy (anonymous):

lcm of 2 numbers is 45 times their hcf. if one of the numbers is 125 and the sum of the hcf and lcm is 1150.find the other no.

OpenStudy (ganpat):

L.C.M. x H.C.F. = 1st number x 2nd number one number is 125, and HCF + LCM = 1150 also, LCM = 45 HCF, HCF + 45HCF = 1150 46 HCF = 1150 So, HCF = 25... SO putting all values in equation.. L.C.M. x H.C.F. = 1st number x 2nd number 45 (25) * 25 = 125 * x so x = 225... is the other number.. is that right ?????